# Object-Centric Concept-Bottlenecks

### v2

@@ 1 Introduction
In this work, we aim to bridge this gap by integrating the advantages of object-centric modeling into the CBM framework, with a focus on models that can handle complex, natural images while maintaining interpretability through structured, object-level representations.
To this end, we introduce the first Object-Centric Concept Bottlenecks (OCB). In this framework, an object proposal module first identifies relevant objects in an image, which are then represented by high-level concepts from a concept discovery module. For the final task prediction, concept activations from both the original image and individual objects are aggregated and passed through a linear prediction layer. Following recent trends in interpretable-by-design research, OCB utilizes pretrained models for both object detection and concept discovery, thereby reducing the amount of required human supervision. As the addition of object-level representations expands the concept space, effectively aggregating these becomes a new challenge, affecting both interpretability and performance. Thus, our evaluations include a detailed analysis of the critical components of OCB. We conduct experiments across multiple benchmark datasets, including our newly introduced COCOLogic benchmark, based on MSCOCO (Lin et al., 2014), which requires models to classify images according to concepts defined by logical combinations of objects. Importantly, we also evaluate CBMs via OCB for the first time on multi-label image classification benchmarks.
Overall, our work shows that Object-Centric Concept Bottlenecks represents an important step towards more competitive, yet interpretable AI models, with a particular focus on analyzing the potential of object-centricity and applying it to CBMs for natural images in the wild. Our contributions can be summarised as:
@@
The remainder of the paper is structured as follows. We begin with a review of related work, highlighting recent developments in the field. We then introduce our proposed framework, Object-Centric Concept Bottlenecks, and present its formal description. This is followed by a comprehensive experimental evaluation, including several ablation studies that examine key components of the framework. Finally, we discuss our findings and conclude the paper.
@@ 2 Related Work
Concept-Based Models. The introduction of Concept Bottleneck Models (CBMs) (Koh et al., 2020) marked an important moment in the growing interest in concept-based models (cf. (Yeh et al., 2021; Fel et al., 2023) for overviews on concept-based explainability). Their appeal lies in the promise of interpretable predictions and a structured interface for human interaction. While the original CBM framework relied on fully supervised concept annotations, subsequent research has relaxed this requirement by introducing unsupervised concepts (Sawada and Nakamura, 2022), leveraging pretrained vision-language models like CLIP for concept extraction (Bhalla et al., 2024; Yang et al., 2023; Oikarinen et al., 2023; Panousis et al., 2024; Yamaguchi et al., 2025), or employing fully unsupervised concept discovery methods (Stammer et al., 2024b; Schrodi et al., 2024; Schut et al., 2025). Efforts have also focused on improving the bottleneck interface by mitigating concept leakage (Havasi et al., 2022) and enabling dynamic expansion of the concept space (Shang et al., 2024). Despite this progress, prior work has not explored the potential of explicitly encoding object-centric concept representations. Although some studies have incorporated such representations into bottleneck models (Yi et al., 2018; Stammer et al., 2021; Wüst et al., 2024; Delfosse et al., 2024), these approaches have thus far been limited to synthetic data. In contrast, our work introduces object-centric bottlenecks for real-world image data, bridging this gap in the literature.
@@
Utilizing Pre-trained Models for Interpretability. A growing body of work investigates how large pretrained models, such as vision-language models (VLMs) and large language models (LLMs), can be leveraged to enhance interpretability, particularly through concept bottlenecks. Many recent approaches incorporate features from pretrained models into inherently interpretable architectures to avoid the need for manual concept supervision. For instance, Yang et al. (2023) and Rao et al. (2024) extract concepts from language models or detection backbones, while Ismail et al. (2024) introduce CB-pLMs for controllable and interpretable protein design. In natural language processing, Tan et al. (2024) retrofit LLMs with lightweight bottlenecks, and Sun et al. (2024) scale this to classification and generation with competitive performance and built-in interpretability. Further, Chen et al. (2025) improve self-explanation capabilities by aligning internal LLM representations with explicit concept subspaces. OCB is aligned with these works and utilizes pretrained models to build object-centric concept representations.
@@ 3 Object-Centric Concept Bottlenecks (OCB)
From Images to Objects. To incorporate object representations into concept-based models, OCB first identifies relevant objects within an image (see (I) in Fig.  2). This is achieved using an object proposal model $o:\mathcal{X}\rightarrow(\mathcal{B},\mathcal{S})$, which maps an input image $x$ to a set of bounding boxes $B$ and associated certainty scores $S$, indicating the confidence that each box contains a meaningful object. Since $o$ may generate numerous proposals of varying quality, OCB filters these object proposals to remove noisy object proposals. Bounding boxes that are too small (low resolution and less likely to carry meaningful concepts) or too large (too similar to the full image) are removed based on configurable size thresholds $t_{min}$ and $t_{max}$. Additionally, boxes with certainty scores below the threshold $t_{cer}$ are discarded. As many object-proposal models have a tendency to generate multiple bounding boxes for the same object, all object proposals with a high Intersection over Union (IoU) $>t_{\mathrm{IoU}}$ are filtered out, inspired by non-maximum suppression (Neubeck and Van Gool, 2006). For this, OCB first orders the object-proposals according to their certainty scores and removes all bounding boxes that have an IoU higher than $t_{\mathrm{IoU}}$ with any remaining bounding boxes of a higher certainty (presented here as $\mathrm{IoU}(b,B)>t_{\mathrm{IoU}}$). Lastly, we restrict the number of object proposals to a maximum of $k$ bounding boxes, which can, for example, be set based on the complexity of the task and the images (cf. full filtering pseudocode in Alg.  1 in the appendix). Based on this filtering process, we obtain a refined set of object proposals:
@@
Together, our enriched concept activations for in input image consist of the activations for the whole image and all object proposals: $c=\{c_{x}\}\cup\{c_{B}\}=\{c_{x},c_{0},\cdots,c_{k}\}\in\mathbb{R}^{(k+1)C}$.
@@
The max aggregation takes the element-wise maximum over the concept-activations: $\texttt{max}(c)^{l}=\operatorname*{arg\,max}_{j\in(x,0,\cdots,k)}c_{j}^{l}$. This maintains the same value range for all activations but loses information about activation strength or number. The sum aggregation maintains the information about the individual activation strengths: $\texttt{sum}(c)^{l}=\sum_{j\in(x,0,\cdots,k)}c_{j}^{l}$. However, this aggregation still does not have explicit information about the number of times a concept has been activated, and thus about counts of objects in the image. The count aggregation is explicitly suited for situations where numbers of objects are important: $\texttt{count}(c)^{l}=\sum_{j\in(x,0,\cdots,k)}\mathbf{I}(c_{j}^{l})$, where $\mathbf{I}(a)$ equals $1$ if $a\neq 0$ and 00 otherwise. Compared to concat, all these aggregations have the advantage that the output space remains $\mathbb{R}^{C}$, independent of the number of objects that concept activations are computed for. However, as all of these aggregations also lose some kind of information, the final aggregation method we introduce is sum + count. This aggregation combines sum and count: $\texttt{sum + count}(c)^{l}=(\texttt{sum}(c)^{l},\texttt{count}(c)^{l}$, with an output space of $\mathbb{R}^{2C}$. This aggregation retains information about the concept activation strength as well as explicitly keeping track of concept activation counts, while still having an output space that is constant in size relative to the number of objects. We provide evaluations on the different forms of aggregations in Sec.  5.
@@ 4 The COCOLogic Dataset
Previous CBM research has primarily focused on simple single-label classification tasks, often using datasets where success depends largely on detecting the presence of a specific object type. In contrast, our work extends the evaluation of CBMs to both more complex multi-label settings and, for the first time, to a novel single-label classification task that demands richer reasoning capabilities on natural images. For this, we introduce COCOLogic, a new benchmark based on the MSCOCO dataset (Lin et al., 2014), that is designed to evaluate a model’s ability to perform complex reasoning over natural visual content in a single-label classification context.
COCOLogic requires models to classify images based on high-level semantic concepts defined by logical combinations of objects. These concepts incorporate disjunctions, conjunctions, negations, and counting constraints over detected object categories, far beyond simple object presence. Each image is assigned one of ten mutually exclusive labels, determined by a logical rule applied to COCO object annotations. Two examples are shown in Fig.  3 (cf. Sec.  D for more details).
The dataset construction process ensures exclusive labeling, such that each image satisfies exactly one class definition. This setting allows for a controlled comparison of model expressiveness on complex visual reasoning tasks, while still formulated in the context of single-label classification. COCOLogic thus serves as a bridge between (logical) reasoning and visual understanding within a simple training setup, providing a dataset that is both grounded in realistic visual scenes and challenging from a logical reasoning perspective.
@@ 5 Experimental Evaluations
Metrics. When comparing downstream performance of the investigated models on the single-label datasets, we follow recent work and report test accuracy. For the multi-label datasets, we report mean Average Precision (mAP). In all experiments, we report average and standard deviation over 5 seeds.
@@
Models. For our evaluations, we do not rely on any concept or object-level supervision but leverage the potential of pretrained models for instantiating the different components of OCB. For the object-proposal model $o$ we explore the use of the "narrow-purpose" pretrained model MaskRCNN (He et al., 2017) (denoted as RCNN), which has been trained to detect objects of the COCO (Lin et al., 2014) dataset. As a second, "general-purpose" proposal model, we revert to SAM (Kirillov et al., 2023), which has been pretained on SA-1B and, in principle, allows to segment arbitrary images. To instantiate the concept discovery module $h$, it is in principle possible to use any of the recent pretrained CBM approaches. We revert to the practical and quite general SpLiCE (Bhalla et al., 2024) approach, in particular due to its task-independent concept vocabulary. Moreover, SpLiCE can be seen as a generalized framework that subsumes a broad family of recent concept bottleneck models that leverage pretrained vision-language representations such as CLIP (Yang et al., 2023; Oikarinen et al., 2023; Panousis et al., 2024; Yamaguchi et al., 2025). This generality allows SpLiCE to serve as a representative instantiation for studying concept-based models in a unified framework. Additionally, the sparsity in SpLiCE’s resulting concept space aids understandability and thus downstream interpretability.
As a baseline, we construct a CBM based on SpLiCE, where we train a linear layer to perform multiclass or multilabel classification based on the SpLiCE encodings of the whole image. We denote this baseline model as CBM. The concept space of SpLiCE is naturally sparse, which is an important aspect in maintaining its interpretability despite the large general concept vocabulary. As OCB generates multiple object proposals and then concept activations for these objects, the resulting concept space of OCB is slightly less sparse than the concept space of the base CBM with SpLiCE. To validate whether this increase in capacity is the reason for the performance improvement, in our evaluations, we investigate a second baseline: CBM (equal capacity). In this setup, we reduced the sparsity regularization of SpLiCE to have an equal capacity (number of non-zero concept activations) to OCB’s observed one.
5.1 Enhancing Task Performance via Objects (RQ1).
To investigate OCB’s general predictive performance, we compare the performance of CBMs with and without objects on the five different datasets. As the results in Tab.  1 show, the performance of OCB is superior to the non-object-centric CBM over all datasets, i.e., both across the multi-label as well as single-label settings (we hereby focus on the OCB (RCNN) results and compare these to OCB (SAM) in the next paragraph and discuss impact of the aggregation method and the value of $k$ in the context of RQ3 and RQ4, respectively). Interestingly, OCB outperforms the base CBM even on SUN397, despite that dataset consisting of many classes like "sky", "desert sand", or "mountain", where an object-centric approach intuitively would not provide many benefits. However, the performance increase is rather small compared to the improvement on the more object-based tasks. The largest improvement can be seen on COCOLogic, which requires object-based visual reasoning. For COCO(l), we also observe a large boost via object representations, indicating that OCB successfully extends the concept space by information about specific objects in the image. Lastly, while the results in Tab.  1 indicate that more capacity generally improves baseline performance, adding this capacity in an object-centric fashion is substantially better than just having more non-zero image-level concepts.
@@
[…]
@@
Figure 4: There is no one-fits-all choice of the aggregation method. Comparing the performance of different aggregation strategies with OCB (RCNN) and $k=7$.
Although max aggregation leads to the best overall performance, the results indicate that there is no one-fits-all choice of the aggregation method. sum and max perform generally well, while sum + count can have the potential for good performance, but with the caveat of more difficult training of the predictor network. Additionally, it is noteworthy that concat retains full traceability between objects and concepts for model explanations, which might outweigh the slight loss in performance depending on the application. We discuss this further in Sec.  6.
Figure 5: Adding more object proposals tends to improve performance. OCB (RCNN) performance for different values of $k$.
5.4 Impact of the Number of Objects (RQ4).
Let us now turn to the influence of the number of objects for OCB. When refining the object proposals, OCB limits the number of objects to the top $k$. We investigate the impact of this parameter on the performance of OCB in Fig.  5, where the best-performing aggregation method is shown for each dataset. In general, we observe a trend indicating that adding more objects improves performance, in particular for PASCAL-VOC and COCO(h). On the other hand, performance declines beyond three objects on COCOLogic. This is likely due to the nature of COCOLogic: This dataset only requires the recognition of few key objects for correct classification while several classes require category counts or explcude certain categories. Thus, additional low-quality proposals, such as faulty or redundant detections, can lead to incorrect predictions. A similar effect occurs on SUN397, where performance slightly decreases with more objects due to the datasets scene-centric nature, where extra object crops add noise rather than informative concepts. In summary, while more object proposals generally improves performance on object-centric datasets, careful filtering is essential to mitigate noise in tasks sensitive to false positives.
@@
In the previous sections, we investigated whether the addition of object-centric concepts to CBMs can increase task performance. However, another important aspect of CBMs is their inherent interpretability via inspections of the concept space. Indeed OCB’s concept space also allows for a more fine-grained and modular analysis of these concepts. As each object-proposal is directly associated with the discovered concepts, it is possible to provide a conceptual description for each found object. Fig.  6 (top) shows an example where the object-centric concepts enrich the conceptual explanation of the image via previously undiscovered concepts, like the presence of a fork or a wine glass. In contrast, Fig.  6 (bottom) highlights that the object-centric concepts do not always have to extend the concept space by additional concepts. Even if the main concept (Pelican) is already identified in the full image encoding, the object-centric explanations reveal that there is not just one, but two pelicans in the image. Overall, the object-centric concept space allows for a more fine-grained representation of the image, also including information like object counts, which is an important basis for object-centric explanations.

### v4

@@ 1 Introduction
In this work, we aim to bridge this gap by integrating the advantages of object-centric modeling into CBMs, with a focus on models that can handle complex, natural images while maintaining interpretability through structured, object-level representations.
To this end, we introduce Object-Centric Concept Bottlenecks (OCB). In this framework, an object proposal module first identifies relevant objects in an image, which are then represented by high-level concepts from a concept discovery module. For the final task prediction, concept activations from both the original image and individual objects are aggregated and passed through a linear prediction layer. Following recent trends in interpretable-by-design research, OCB utilizes pretrained models for both object detection and concept discovery, thereby reducing the amount of required human supervision. As the addition of object-level representations expands the concept space, effectively aggregating these becomes a new challenge. Thus, our evaluations include a detailed analysis of the critical components of OCB. We conduct experiments across multiple datasets, including our newly introduced COCOLogic benchmark, based on MSCOCO (Lin et al., 2014), which requires models to classify images according to concepts defined by logical combinations of objects. Importantly, we also evaluate CBMs via OCB for the first time on multi-label image classification benchmarks.
Overall, our work shows that Object-Centric Concept Bottlenecks represents an important step towards more competitive, yet interpretable AI models, with a particular focus on object-centricity and its application to CBMs for natural images. Our contributions can be summarised as:
@@
The remainder of the paper is structured as follows. We begin with a review of related work, highlighting recent developments in the field. We then introduce our proposed framework, Object-Centric Concept Bottlenecks, and present its formal description. This is followed by a comprehensive experimental evaluation, including several ablation studies that examine key components of the framework. Finally, we discuss our findings and conclude the paper. [^1] Code and data available at: https://github.com/DavSte13/Object-Centric-Concept-Bottlenecks/
@@ 2 Related Work
Concept-Based Models. The introduction of Concept Bottleneck Models (CBMs) (Koh et al., 2020) marked an important moment in the growing interest in concept-based models (cf. (Yeh et al., 2021; Fel et al., 2023) for overviews on concept-based explainability). Their appeal lies in the promise of interpretable predictions and a structured interface for human interaction. While the original CBM framework relied on fully supervised concept annotations, subsequent research has relaxed this requirement by introducing unsupervised concepts (Sawada and Nakamura, 2022), leveraging pretrained vision-language models like CLIP for concept extraction (Bhalla et al., 2024; Yang et al., 2023; Oikarinen et al., 2023; Panousis et al., 2024; Yamaguchi et al., 2025), or employing fully unsupervised concept discovery methods (Stammer et al., 2024b; Schrodi et al., 2024; Schut et al., 2025). Efforts have also focused on improving the bottleneck interface by mitigating concept leakage (Havasi et al., 2022) and enabling dynamic expansion of the concept space (Shang et al., 2024). Despite this progress, prior work has not explored the potential of explicit object-centric representations. Although some studies have incorporated such representations into bottleneck models (Yi et al., 2018; Stammer et al., 2021; Wüst et al., 2024; Delfosse et al., 2024), these approaches have thus far been limited to synthetic data. In contrast, our work introduces object-centric bottlenecks for real-world image data. While Prasse et al. (2025) utilizes foundation models to extract an object-based concept bank from images, they compute a single concept activation vector per image, in contrast to OCB, which explicitly combines image and object-level concept activations.
@@
Utilizing Pre-trained Models for Interpretability. A growing body of work investigates how large pretrained models, such as vision-language models (VLMs) and large language models (LLMs), can be leveraged to enhance interpretability, particularly through concept bottlenecks. Many recent approaches incorporate features from pretrained models into inherently interpretable architectures to avoid the need for manual concept supervision. For instance, Yang et al. (2023) and Rao et al. (2024) extract concepts from language models or detection backbones, while Ismail et al. (2024) introduce CB-pLMs for controllable protein design. In natural language processing, Tan et al. (2024) retrofit LLMs with lightweight bottlenecks, and Sun et al. (2024) scale this to classification and generation with competitive performance and built-in interpretability. Further, Chen et al. (2025) improve self-explanations by aligning internal LLM representations with explicit concept subspaces. In line with these works, OCB utilizes pretrained models to build object-centric concept representations.
@@ 3 Object-Centric Concept Bottlenecks (OCB)
From Images to Objects. To incorporate object representations into concept-based models, OCB first identifies relevant objects within an image (see (I) in Fig.  2). This is achieved using an object proposal model $o:\mathcal{X}\rightarrow(\mathcal{B},\mathcal{S})$, which maps an input image $x$ to a set of bounding boxes $B$ and associated certainty scores $S$, indicating the confidence that each box contains a meaningful object. Since $o$ may generate numerous proposals of varying quality, OCB filters these object proposals to remove noisy object proposals. Bounding boxes that are too small (low resolution and less likely to carry meaningful concepts) or too large (too similar to the full image) are removed based on configurable size thresholds $t_{min}$ and $t_{max}$. Additionally, boxes with certainty scores below the threshold $t_{cer}$ are discarded. As many object-proposal models have a tendency to generate multiple bounding boxes for the same object, all object proposals with a high Intersection over Union (IoU) $t_{\mathrm{IoU}}$ are filtered out, inspired by non-maximum suppression (Neubeck and Van Gool, 2006). For this, OCB first orders the object-proposals according to their certainty scores and removes all bounding boxes that have an IoU higher than $t_{\mathrm{IoU}}$ with any remaining bounding boxes of a higher certainty (presented here as $\mathrm{IoU}(b,B)>t_{\mathrm{IoU}}$). Lastly, we restrict the number of object proposals to a maximum of $k$ bounding boxes, which can, for example, be set based on the complexity of the task and the images (cf. full filtering pseudocode in Alg.  1 in the appendix). Based on this filtering process, we obtain a refined set of object proposals:
@@
Together, our enriched concept activations for an input image consist of the activations for the whole image and all object proposals: $c=\{c_{x}\}\cup\{c_{B}\}=\{c_{x},c_{0},\cdots,c_{k}\}\in\mathbb{R}^{(k+1)C}$. By default, the concept space for image-level concepts and object-level concepts is the same. If this is the case, the underlying assumption of OCB is that these concept spaces are sparsely activated, so that not every concept is active (to some extent) for every image and object.
@@
The max aggregation takes the element-wise maximum over the concept-activations: $\texttt{max}(c)^{l}=\operatorname*{arg\,max}_{j\in(x,0,\cdots,k)}c_{j}^{l}$. This maintains the same value range for all activations but loses information about activation strength or number. The sum aggregation maintains the information about the individual activation strengths: $\texttt{sum}(c)^{l}=\sum_{j\in(x,0,\cdots,k)}c_{j}^{l}$. However, this aggregation still does not have explicit information about the number of times a concept has been activated, and thus about counts of objects in the image. The count aggregation is explicitly suited for situations where numbers of objects are important: $\texttt{count}(c)^{l}=\sum_{j\in(x,0,\cdots,k)}\mathbf{I}(c_{j}^{l})$, where $\mathbf{I}(a)$ equals $1$ if $a\neq 0$ and 0 otherwise. Compared to concat, all these aggregations have the advantage that the output space remains $\mathbb{R}^{C}$, independent of the number of objects that concept activations are computed for. However, as all of these aggregations also lose some kind of information, the final aggregation method we introduce is sum + count. This aggregation combines sum and count: $\texttt{sum + count}(c)^{l}=(\texttt{sum}(c)^{l},\texttt{count}(c)^{l}$, with an output space of $\mathbb{R}^{2C}$. This aggregation retains information about the concept activation strength as well as explicitly keeping track of concept activation counts, while still having an output space that is constant in size relative to the number of objects. We provide evaluations on the different forms of aggregations in Sec.  5.
@@ 4 The COCOLogic Dataset
Previous CBM research has primarily focused on simple single-label classification tasks, often using datasets where success depends largely on detecting the presence of a specific object type. In contrast, our work extends the evaluation of CBMs to both more complex multi-label settings and, for the first time, to a novel single-label classification task that demands richer reasoning capabilities on natural images. For this, we introduce COCOLogic, a new benchmark based on the MSCOCO dataset (Lin et al., 2014), that is designed to evaluate a model’s ability to perform structured visual reasoning in a single-label classification context.
COCOLogic is intended as a proxy task for structured reasoning in vision, capturing aspects such as multi-object composition, negation, and hierarchical categories while still being grounded in real-world data. Concretely, COCOLogic requires models to classify images based on high-level semantic concepts defined by logical combinations of objects. These concepts incorporate disjunctions, conjunctions, negations, and counting constraints over detected object categories, far beyond simple object presence. Each image is assigned one of ten mutually exclusive labels, determined by a logical rule applied to COCO object annotations. Two examples are shown in Fig.  3 (cf. Sec.  D for more details).
The dataset construction process ensures exclusive labeling, such that each image satisfies exactly one class definition. This setting allows for a controlled comparison of model expressiveness on complex visual reasoning tasks, while still formulated in the context of single-label classification. COCOLogic thus serves as a bridge between logical reasoning and visual understanding within a simple training setup: it remains grounded in realistic visual scenes while being inherently challenging from a logical reasoning perspective.
@@ 5 Experimental Evaluations
Metrics. When comparing downstream performance of the investigated models on the single-label datasets, we follow recent work and report test accuracy (balanced accuracy for COCOLogic due to its strong class imbalance). For the multi-label datasets, we report mean Average Precision (mAP). In all experiments, we report average and standard deviation over 5 seeds.
@@
Models. For our evaluations, we do not rely on any concept or object-level supervision but leverage the potential of pretrained models for instantiating the different components of OCB. For the object-proposal model $o$ we explore the use of the "narrow-purpose" pretrained model MaskRCNN (He et al., 2017) (denoted as RCNN), which has been trained to detect objects of the COCO (Lin et al., 2014) dataset. As a second, "general-purpose" proposal model, we revert to SAM (Kirillov et al., 2023), which has been pretained on SA-1B and, in principle, allows to segment arbitrary images. To instantiate the concept discovery module $h$, it is in principle possible to use any of the recent pretrained CBM approaches. We revert to the practical and quite general SpLiCE (Bhalla et al., 2024) approach, in particular due to its task-independent concept vocabulary. Moreover, SpLiCE can be seen as a generalized framework that subsumes a broad family of recent concept bottleneck models that leverage pretrained vision-language representations such as CLIP (Yang et al., 2023; Oikarinen et al., 2023; Panousis et al., 2024; Yamaguchi et al., 2025). This generality allows SpLiCE to serve as a representative instantiation for studying concept-based models in a unified framework. Additionally, SpLiCE generates sparse concept activations, which aid understandability and thus downstream interpretability.
We first evaluate a non-interpretable model that maps CLIP image embeddings to class predictions via a two-layer MLP, serving as an upper bound on performance using CLIP features alone. For interpretable baselines, we use a SpLiCE-based CBM, where a linear classifier predicts multiclass or multilabel outputs from whole-image SpLiCE encodings. SpLiCE’s inherently sparse concept space aids interpretability despite its large vocabulary. Our method (OCB) generates concepts from multiple object proposals, yielding a slightly less sparse concept space than the base CBM. To test if this increased capacity drives performance gains, we include a CBM (equal capacity) variant with reduced SpLiCE sparsity regularization to match OCB ’s number of active concepts. We also compare to the Coarse-to-Fine CBM (C2F-CBM) (Panousis et al., 2024), which uses hierarchical, patch-based concepts. For fairness, we adopt the same LAION-based vocabulary as OCB and disable OCB sparsity constraints. Since C2F-CBM supports only single-label classification, it is evaluated on SUN397 and COCOLogic. Lastly, we compare to an opaque model that maps CLIP image embeddings directly to class predictions via a two-layer MLP (CLIP+MLP).
@@
CLIP+MLP $89.60\mbox{\scriptsize$\pm$ 0.05}$ $89.67\mbox{\scriptsize$\pm$ 0.03}$ $68.20\mbox{\scriptsize$\pm$ 0.13}$ $79.62\mbox{\scriptsize$\pm$ 0.11}$ $65.95\mbox{\scriptsize$\pm$ 1.00}$
@@
CBM (equal capacity) $82.90\mbox{\scriptsize$\pm$ 0.01}$ $86.31\mbox{\scriptsize$\pm$ 0.00}$ $61.70\mbox{\scriptsize$\pm$ 0.00}$ $74.85\mbox{\scriptsize$\pm$ 0.01}$ $60.01\mbox{\scriptsize$\pm$ 0.09}$
@@
5.1 Enhancing Task Performance via Objects (RQ1).
To investigate OCB’s general predictive performance, we compare the performance of CBMs with and without objects on the five different datasets. As the results in Tab.  1 show, the performance of OCB is superior to the non-object-centric CBM over all datasets, i.e., both across the multi-label as well as single-label settings (we hereby focus on the OCB (RCNN) results and compare these to OCB (SAM) in the next subsection). Interestingly, OCB outperforms the base CBM even on SUN397, despite that dataset consisting of many classes like "sky", "desert sand", or "mountain", where an object-centric approach intuitively would not provide many benefits. However, the performance increase is rather small compared to the improvement on the more object-based tasks. The largest improvement can be seen on COCOLogic, which requires object-based visual reasoning. For COCO(l), we also observe a large boost via object representations, indicating that OCB successfully extends the concept space by information about specific objects in the image. To better contextualize our results, we also compare against a non-interpretable upper bound that uses the same image backbone as our models (CLIP+MLP). We observe that CLIP+MLP achieves higher accuracy, particularly than vanilla concept-based approaches, at the cost of offering no meaningful explanations for its predictions. However, our object-centric CBM (OCB) substantially narrows this gap: by leveraging structured object-level concept representations, OCB recovers much of the predictive power of the opaque model while retaining transparency.
As Tab.  1 indicates that more concept capacity improves the baseline performance, we compare base CBM and OCB with different concept capacities (i.e., the average number of non-zero concepts, cf. Fig.  8 in the appendix). Overall, increasing concept capacity improves the performance of the baseline and OCB, but on all object-based tasks, adding object-level concepts is consistently superior to adding more image-level concepts, again indicating that OCB allows for more competitive CBMs.
Table 2: Comparison between patch vs object encodings. Results for SUN397 and COCOLogic datasets. Best results are in bold.
Model SUN397 COCOLogic
C2F-CBM $70.12\mbox{\scriptsize$\pm$ 0.24}$ $65.81\mbox{\scriptsize$\pm$ 1.57}$
OCB (non-sparse) $\mathbf{89.60}\mbox{\scriptsize$\pm$ 0.57}$ $\mathbf{68.63}\mbox{\scriptsize$\pm$ 0.42}$
Lastly, in Tab.  2 we evaluate against the recent Coarse-to-Fine CBM (Panousis et al., 2024), which introduces hierarchical, patch-based image processing. As C2F does not support multi-label classification, we perform these evaluations on the SUN397 and COCOLogic datasets and, for fair comparisons, utilise a common LAION-based concept vocabulary with no sparsity constraints. We observe that our OCB model consistently outperforms C2F under these conditions, highlighting the benefit of object-centric representations over patch-based hierarchies.
@@
Figure 4: Different datasets benefit from different aggregation methods. While the performance comparison of different aggregation strategies with OCB (RCNN) and $k=7$ does not show a one-fits-all choice, max or sum are always solid choices.
Although max aggregation leads to the best overall performance, the results indicate that there is no one-fits-all choice of the aggregation method. sum and max perform generally well, while sum + count can have the potential for good performance, but with the caveat of more difficult training of the predictor network. Additionally, it is noteworthy that concat retains full traceability between objects and concepts for model explanations, which might outweigh the smaller performance gain depending on the application (cf. discussion in Sec.  6).
5.4 Impact of the Object Proposal Refinement (RQ4).
Let us now turn to the influence of the number of object proposals for OCB. The most influential element of the refinement is the limitation of the number of objects to the top $k$. We investigate the impact of this parameter on the performance of OCB in Fig.  5, where the best-performing aggregation method is shown for each dataset. In general, we observe a trend indicating that adding more objects improves performance, in particular for PASCAL-VOC and COCO(h). On the other hand, performance beyond a certain point degrades again for COCOLogic and SUN397.
Figure 5: Adding more object proposals improves performance on object-based tasks. However, for complex tasks like COCOLogic, adding to many (noisy) object proposals can result in reduced performance. OCB (RCNN) performance for different values of $k$.
Due to the structure of COCOLogic, only the recognition of few key objects is necessary for a correct classification, while several classes require category counts or exclude certain categories. Thus, additional low-quality proposals, such as faulty or redundant detections, can lead to incorrect predictions. A similar effect occurs on the scene-centric SUN397, where extra object crops add noise rather than informative concepts. We also investigate the effect $t_{\min}$, the threshold limiting the minimum object size in Tab.  10 in the appendix. Here, it can be observed that a lower threshold generally increases performance, with the exception of SUN397, where the scene-centric nature of this dataset does not benefit from the detection of small objects. In summary, allowing more object proposals and the detection of smaller objects generally improves on object-centric datasets, careful filtering is essential to mitigate noise in tasks sensitive to false positives.
@@
In the previous sections, we investigated whether the addition of object-centric concepts to CBMs can increase task performance. However, another important aspect of CBMs is their inherent interpretability via inspections of the concept space. Indeed OCB’s concept space also allows for a more fine-grained and modular analysis of these concepts. As each object proposal is directly associated with the discovered concepts, it is possible to provide a conceptual description for each found object. Fig.  6 (top) shows an example where the object-centric concepts enrich the conceptual explanation of the image via previously undiscovered concepts, like the presence of a fork or a wine glass. In contrast, Fig.  6 (bottom) highlights that the object-centric concepts do not always have to extend the concept space by additional concepts. Even if the main concept (Pelican) is already identified in the full image encoding, the object-centric explanations reveal that there is not just one, but two pelicans in the image. Overall, the object-centric concept space allows for a more fine-grained representation of the image, also including information like object counts, which is an important basis for object-centric explanations.